OpenStudy (anonymous):

if y varies directly as y and x =1 as y = 4 then what is the constant of variation

OpenStudy (kainui):

So if they vary directly, then that means you can multiply one number by the other to get to the other number. And in this case they say that it's a constant. So what number can you multiply x by to get y?

OpenStudy (anonymous):

constant of variation would be = y/x = 4/1 =4

OpenStudy (anonymous):

Thanks all !